OpenStudy (anonymous):

write the slope intercept form of the equation of the line (6, -6), (-6, -4)

OpenStudy (anonymous):

(x1,y1) = (6,-6) (x2,y2) = (-6,-4) 1) calculate the slope m = (y2-y1)/(x2-x1) 2) calculate the y-intercept,b, y = m*x +b you may insert either one of the two points you have above for the values of y and x 3) take the values of slope and y-intercept and substitute them for m and b in the equation above. 4) slope-intercept form: y = m*x + b

OpenStudy (anonymous):

ok may you let me know if i am right when i am done

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

-2/12 is my slope

OpenStudy (anonymous):

yep you can reduce it to -1/6

OpenStudy (anonymous):

do i have to reduce it or would i get the same answer

OpenStudy (anonymous):

you will get the same answer, reducing just helps make the calculations easier

OpenStudy (anonymous):

b= -7

OpenStudy (anonymous):

ah it will be -5, bc you added a +1 to a -6 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I took (6,-6) as the (x,y) values, and m as -1/6 for easier calculations

OpenStudy (anonymous):

\[-6 = (-1/6)*(6) + b\]\[-6 = -1 + b\]\[b = -5\]

OpenStudy (anonymous):

i dont understand step 3

OpenStudy (anonymous):

step 3 was just to get your completed answer, now that you have the value for m,slope, and b, y-intercept put them back into the equation y = m*x +b, leaving y and x as variables

OpenStudy (anonymous):

that is slope-intercept form

OpenStudy (anonymous):

y=-2/12x+(-5)

OpenStudy (anonymous):

yep that is correct

OpenStudy (anonymous):

good luck on the rest of your problems

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome hon